OpenStudy (anonymous):

Solve for r in the equation C = 2d - 2r.

OpenStudy (saifoo.khan):

\[r = -\frac{C - 2d}{2} \]

OpenStudy (anonymous):

is the negetive sign in fornt of the c or 2?

OpenStudy (saifoo.khan):

it's before the whole fraction.

OpenStudy (anonymous):

A) r = C + 2d B) r = (-C+2d)/2 C) r = C + 2d - 2 D) r = (C+2d)/-2